OpenStudy (anonymous):

Write the equation in standard form of the circle whose center is at (-4,1) that has radius 10 units.

OpenStudy (anonymous):

Do you know the standard form for the equation of a circle?

myininaya (myininaya):

the standard form is (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

\[(x-h)^2 + (y - k)^2 = r^2\] Where (h,k) is the center point of the circle, and r is the radius.